OpenStudy (anonymous):

add and simplify and show work 7 3 --- + --- x+6 x²+6x

OpenStudy (anonymous):

numerators- top of the fraction can only be added if the denominators are the same

geerky42 (geerky42):

\[\dfrac{7}{x+6} + \dfrac{3}{x^2 + 6x} = \dfrac{7x}{x(x+6)} + \dfrac{3}{x^2 + 6x} \\ = \dfrac{7x+3}{x^2 + 6x}\]